[Halld-jlab] Hall D Staff Meeting Notes, July 14, 2014

Mark M. Ito marki at jlab.org
Mon Jul 14 18:12:16 EDT 2014


Folks,

Find the notes below and at 
https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_14,_2014 
.

   -- Mark
_____________________

Private:Hall D Staff Meeting, July 14, 2014

    Present: Alex Barnes, Eugene Chudakov (chair), Mark Dalton, Alexandre
    Deur, Sergey Furletov, Mark Ito, David Lawrence, Lubomir Pentchev,
    Elton Smith, Sascha Somov, Simon Taylor, Glenn Young

Eugene

      * various plots of progress were shown
           + there are three months left for construction, 1.1 $M left
           + pre-ops will go until December, 120 k$ left of original 180 k$
           + construction categories:
                o online will not catch up, need to prioritize, may need a
                  special meeting
                o target overspending by quite a bit
                o components behind
                o infrastructure looks OK
           + pre-ops categories (overall a bit late)
                o solenoid spending overspent
                o online and software checkout behind
                o beamline behind
      * post-doc search has started
           + committee: Eugene, Elton, Glenn, Lubomir, Mark Jones
           + no short list yet
      * DIRC
           + Rolf willing to spend on R&D for this fiscal year
           + new post-doc may work on this at some level

Sergey F.

      * working on start-up of large-scale DAQ
      * stand-alone set-up are running as well
      * some crashes, notably on event builder node
           + debugging scheme being instituted by Paul Letta
           + for a full system of 25 crates this can happen after 20
             minutes
      * changing start-up of DAQ
           + include more components
           + COOL set-up space configuration
      * debugging histogram package
           + runs in secondary read-out list
      * re-arranging directory structure for support files
           + e. g., pedestals for flash
           + TOF first system using it

Sascha

      * time-of-flight running with self-trigger through CTP
           + see correlated hits in paddles, both flash and TDC
           + Sasha Ostrovidov and Aristeidis Tsaris are here from FSU
             taking data
      * configuration files for most systems in place, need to check them
      * started reading F1TDCs for BCAL, 6 crates
           + discriminator crates not initializing properly
           + there may be version difference problems with F1TDC
      * started checking GTP to get ready for BCAL cosmic trigger
           + components exist, need to put them together
      * pair spectrometer, course-grained, installed in Hall, connections
        checked
           + next: read-out
      * pair spectrometer, fine-grained, boxes moved to Hall, cable trays
        going in
      * tagger hodoscope
           + Franz here last week, checked counters with sources in EEL 126
           + they had been check at CUA, about 10% show intermittent
             connection when cables wiggled
           + all sections in tagger hall now, connected to patch panel
           + HV not connect to panel yet
      * microscope
           + all bundles fabricated
           + beginning of August delivery

Alex B.

      * here to work on microscope installation
           + delivery expected second full week of August
           + fiber testing going on at UConn with laser diode
           + cosmic testing as well
      * will also work on active collimator
           + instrumentation halfway there
           + all cables are in hand

Lubomir

      * taking FDC data with CDC trigger
           + analyzing one 48-hour run with 14 crates
           + would like to use BCAL trigger for higher occupancy per
             trigger
           + see a few tens of bad channels, mostly in third package
                o some are due to flash
                o one flash is noisy
      * gas system issues being worked through
      * noise issue is being studied, Northwestern grad student on it
           + seen mainly in first and second packages
           + see problem seen by Beni for CDC
           + VME crate switching power supplies a contributor as well
      * CDC: Mike Staib working on noise investigation
           + Beni analyzing data for CDC also

Mark I.

      * Start counter scintillator
           + scratches largely removed with polishing
           + base time resolution before polishing not too far off from
             prototypes
           + with polishing time resolution goes back to good prototype
             level
           + will delay shipment to JLab, early Sep. perhaps
      * will be starting work on data challenge 3
           + use EVIO data in tape library as starting point
           + simulate full-scale reconstruction pass

Simon

      * CDC alignment tests started
           + working as some level with real cosmic data using DAQ plug-in
             for CDC
      * for FDC, using bggen events
      * working on estimate of how much data is needed
           + to do rotation of FDC wire planes, 100 k straight track events
             needed
           + for CDC need 3000 k cosmics
      * working on improving efficiency for multi-track events at track
        candidate state

Mark D.

      * working on BCAL calibration using CDC tracks
           + use track to correct for angle
      * TDCs not analyzed yet

Elton

      * BCAL
           + reading out flash data only
                o systematics seen, e. g., found jumpers not set correctly
                  on flash
                o one confusing channel with half normal gain
                o otherwise gains show 10% spread
           + TDC check-out next
           + Will McGinley[?] working on a mu-to-e trigger
      * FCAL
           + have abandoned AC system in favor of blower
           + Chris Stanislav finishing boxes to replace temperature and
             humidity monitoring system
                o spares can be placed around the Hall
                o use PLC to read out voltage levels
      * RF system
           +
                o Plan A: use coax cable
                o Plan B: use optical fiber
                o Plan C: use standalone 500 MHz clock, modulate with BCM
                  signal
           + Plan B costs 100 k$
           + Plan C does not have detail design yet
           + Alam Toro had electronics design for Plan A, working to finish
             it

Alexandre

      * 6 ion chambers planed for tagger hall
           + will serve as input to FSD system
           + two at entrance to tagger hall
           + two along beam line to dump
           + one near dump
           + one at exit of tagger hall
      * neutron detectors planned for tagger hall and Hall D
           + will borrow from RadCon to start\
      * will also place OSIs in various locations
      * changing design of amorphous radiation to provide FSD signal

David

      * code in JANA now for CAEN TDC
           + almost done, Aristeidis will finish
      * fixed memory leak in DAQ plug-in
      * working bench-marking Hazwell-chip based evaluation computer
           + being considered for purchase for the farm
           + goes back Wednesday
      * DAQ testing sessions scheduled for Tue. and Thu.

Readiness Reviews, Elton and Eugene

      * Aug. 8: review of photon beamline ACC and Hall D
      * Readiness review from two weeks ago:
           + no recommendations
           + comments about more detail needed on commissioning plan
      * ARR[?] review August 26th

-- 
Mark M. Ito, Jefferson Lab, marki at jlab.org, (757)269-5295



More information about the Halld-jlab mailing list